OpenStudy (anonymous):

solve 2(x-6)+4(x+2)-6 I need help

OpenStudy (anonymous):

what are we solving for?x?

OpenStudy (anonymous):

2x-6+4x+8-6 Go From There .

OpenStudy (anonymous):

all the numbers

OpenStudy (anonymous):

the solution

OpenStudy (anonymous):

i have all real numbers

OpenStudy (anonymous):

someone is lazy lol 2x+4x -6+8-6= -4 6x-4 6x=4 4/6=x

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ty